### v1

@@ 5.3 SCALABILITY
As the LC-classes are derived from the power sets of the sample labels, the number of LC-Protonets increases significantly as the number of classes $N$ grows. This results in a corresponding increase in inference time, as the distances from all LC-Protonets must be computed for each query item. In Figure 4, the number of classes $N$ is shown on the $x$-axis, while the left $y$-axis represents the number of LC-Protonets, and the right $y$-axis shows the inference time per query item (in milliseconds), averaged across all datasets.
When $N$ increases from $5$ to $15$, the number of LC-Protonets rises by a factor of 30, from $21$ to $644$, while inference time increases from $16$ to $31$ milliseconds. As the number of classes continues to grow, the number of LC-Protonets increases substantially, reaching $38,953$ for $60$ tags, and the inference time also rises to $912$ milliseconds. However, the gradient of the increase in inference time is less steep compared to the growth in LC-Protonets. We believe these scalability issues can be mitigated by incorporating optimization techniques, such as skipping the computation of distances from LC-Protonets with identical representations that correspond to different LC-classes (which may be hierarchically related). We plan to explore these methods in future work.
It is important to note that the inference process only runs during testing and not during model training. The average training time across all three methods was less than an hour, with no significant differences between them. Finally, the trainable parameters amount to $3.66$ million when training from scratch or performing full fine-tuning, and $262,000$ when only fine-tuning the last layer. The experiments were conducted on an NVIDIA RTX A5000 GPU.

### v2

@@ 5.3 SCALABILITY
As the LC-classes are derived from the power sets of the sample labels, the number of LC-Prototypes increases significantly as the number of classes $N$ grows. This results in a corresponding increase in inference time, as the distances from all LCPs must be computed for each query item. In Figure 4, the number of classes $N$ is shown on the $x$-axis, while the left $y$-axis represents the number of LC-Prototypes, and the right $y$-axis shows the inference time per query item (in milliseconds), averaged across all datasets.
When $N$ increases from $5$ to $15$, the number of LCPs rises by a factor of 30, from $21$ to $644$, while inference time increases from $16$ to $31$ milliseconds. As the number of classes continues to grow, the number of LCPs increases substantially, reaching $38,953$ for $60$ tags, and the inference time also rises to $912$ milliseconds. However, the gradient of the increase in inference time is less steep compared to the growth in LCPs. We believe these scalability issues can be mitigated by incorporating optimization techniques, such as skipping the computation of distances from LCPs with identical representations that correspond to different LC-classes (which may be hierarchically related). We plan to explore these methods in future work.
The inference process runs only during testing and not during model training. The average training time across all three methods was under an hour, with no significant differences between them. The trainable parameters amount to $3.66$ million for training from scratch or full fine-tuning, and $262,000$ for fine-tuning only the last layer. The experiments were conducted on an NVIDIA RTX A5000 GPU.